Tech Staff Meeting Notes July 5, 2017

The top priority is always repair of down stations.
NM problem stations
ET problem stations

Duty Table updated June 2014


News & Highlights

Mitch sent a draft to Chuck for the signpost outside the middle house. Hearing no objections, Chris will proceed.

Development continues between VLF and Francis on a 16-bit raspberry pi based ISIS sensor system with timing (DISIS?). Currently its doing file transfers but streaming, preferably via seedlink or similar, is needed.

Eric's visit was successful and he has what he needs to get SLU loaded into SIS. Mitch learned at the NIC meeting that there is no official AQMS tool to load extened station xml into AQMS. Caltech has a tool that loads inventory and response but the inventory tables are not part of the archdb or rtdb. The response part uses dataless and they are seeing about stripping that out for people to use in the interim (or we could use our old klunky, buggy, tool).

PNSN developed an interface for the IRIS QC tool Mustang ( Its called mustangular and is on github


SIS is fully operational for Memphis stations. Now each person needs to keep it up to date for equipment for which they are responsible (from the June 2014 Duty Table):

Make sure Mitch gets an email for any equipment change that will generate a SIS update. Philip Crotwell has an app that let's you do updates on your phone.


PC's at mttn and hpnc have "puffy caps" and need to be replaced.


A very large list this month.
Data taken from: /dat1/seisadm/Working/NM/201707/20170701_180700_MAN

Misc Problems
BOAR - spikes see image
EPRM - " see image
BETM - " see image
BRNM - " see image
CPAR - "
LVAR - "
TMAR - "
TYAR - "
NAIT.HNZ - see image
RVAR.EHE - see image
HENM.HN* -drifting
HDAR2.HHN.NM.20 bad sensor, see image

Misc AG Problems
X204.HNE see image

Reinstalled DLAR.

Low solar dropouts have several different sources though common problem is an extraordinarily overcast January. Other factors include mis-wiring, ISIS3 xmiters draw more power (so need more solar), AGM's have problems with low power charge, etc. Jim's spreadsheet shows 9 stations that need an upgrade, 10 on the edge, and 13 that need to be checked.
Need to accomplish this task as we visit each station for the annual PM.

Need a site for MORT2.

There was an email discussion on the history of OXF which led to digging up the MOU from 1993. Mitch checked with Dave Wilson at ASL and it may be possible to replace the MOU, which is endorsed at a fairly high level within USGS, with a "station agreement" which is less formal.

I40 Bridge comms.
Received radios for some links. Need to reestablish contacts with law school and get that link going. Need to get a firewall for aztnworm.

Message from Jim Smith re: bridge comm gear from NSMP:

Just letting you know I've ordered all the stuff, and am going to get a quote from AFAR for the Radios. Once I get the Ubiquitis and antennas you listed I'll ship off to you. Should go out by end of next week. Radios will take a while and may come in groups .

The spectre routers have a relatively short character limit for url's. But when you change a config using the web interface with the obsidians, they try to resend the entire configuration using a url and its gets truncated when coupled with a spectre router. So make sure the configs are correct because it requires a site visit to change them.

NODE Tasks

All nodes are due a firmware upgrade that will sufficiently change things such that the the initial upgrade and configure should be done at CERI, then swap hardware at the node. This will be done as machines are replaced.



Replace siding on equipment hut, ASAP.

Put an iboot at MKTA?

Getting interference at the node causing spikes in CPAR, LVAR, LRAR, and TYAR. Receivers have not been aligned for 20 years.



Add a firewall rule to allow local access to tower hardware.

AZTN/HDBR: Need to pull the 200' sensor at the weigh station. Probable water damage.


Roof repair, missing skylights, on the lean-to.

Flooding in the expansion space has become a problem. A Workorder was submitted and P-plant engineers were on-site to assess the problem. The most likely solution will be a retention wall and removal of material to make the foundation above grave. There are several open work orders. Jim will make a prioritized list and speak with zone supervisor.


CVVA needs a new flash card. Steve needs another basalt for a day to configure the new card. NAIT has a bad channel. We'll get a replacement for NAIT, and use the the old NAIT hardware to configure the flash, then send the card to John for installation at CVVA.

Replacement station for GTTN and RBNC being evaluated and permissioned for Southern Highlands Science Learning Center at Purchase Knob (MaggieValley), a part of the Great Smoky Mountains National Park. After installing the compressor isolation hardware the noise study showed no noise from it, but the site did not have hard rock and the signal was very poor. John will send coordinates to Mitch to re-evaluate whether this station can be located elsewhere.


Data taken from: /dat1/seisadm/Working/ET/201707/20170703_015133_MAN


Misc Problems
CPCT.HN* see image
DYTN.HN* -dropping out -as of this morning's email it's been down for over 45 minutes.
CPRT - spikes both short period and strong motion station see image
VHTN - occasional interference - see image

DYTN has regular tick marks. Reftek problem?

CPCT had the RT147 replaced but still has problems. Got the cable for the sm sensor. Waiting on the bb cable to do both at the same time.

GMG has had a power upgrade, new batteries, new solar contoller, and better surge suppression. And it still went down on 1/18. James can get to the radio at the station which is on a separate power feed than the reftek. Jim will send a new BC10 and John will investigate.

John's updated todo list .


Improve RCGA AC power and perform maintenance work on shed.

Move PMTN due to excessive noise and vandalism. The trees at Lone Mountian are too tall to get a signal out. Looking for an alternate site.


GFM2 site is no good. John has ideas for an alternative.


WMTN pit needs to be moved due to erosion.

CPRT comm reconfig. John says, "CPRT is going through the Senior Center at Corryton (cort). We removed everything from svfd an everything but what is on the tower at gttn. cort is indeed using dynamic addressing but data is flowing. Greg Hill the director at the senior center is supposed to be contacting his IT person to see what can be done. I left a message for Greg last week but still haven't heard back. I want to get Greg's IT guy in direct contact with either Steve or James so they can figure it all out."


CVMS has intermittent connection problems with wireless. Its hardwired for now but the wireless needs to be replaced to preserve electical isolation between the building and vault.

AG Network

New AG contract is in negotiation, it may take a couple months. Curb AG spending until then.

Sensor at X40A was replaced.

Vault at Z41A came loose at the bottom--reattached.


The fdl is getting new AC.

The FDL wiki is here:


The seisnet group on umdrive is at You will need your uuid and password.

The seisnet wiki is at You will need your uuid and password.